Mark received 10 ISOs (each option gives him the right to purchase 12 shares of Hendricks Corporation stock for $7 per share) at the time he started working for Hendricks Corporation five years ago when Hendricks’s stock price was $5 per share. Now that Hendricks’s share price is $35 per share, Mark intends to exercise all of his options and hold all of his shares for more than one year. Assume that more than a year after exercise, Mark sells the stock for $35 a share. (Enter all amounts as positive values. Leave no answers blank. Enter zero if applicable.)

a. What are Mark’s taxes due on the grant date, the exercise date, and the date he sells the shares, assuming his ordinary marginal rate is 32 percent and his long-term capital gains rate is 15 percent?

Taxes Due

Grant date _________

Exercise date _________

Sale date _________

b. What are Hendricks’s tax consequences on the grant date, the exercise date, and the date Mark sells the shares, assuming its marginal tax rate is 21 percent?

Taxes Due

Grant date _________

Exercise date _________

Sale date _________
Answer #1

Part A

Grant date = (7*12*10)*32% = $268.80 ($269 if answer required to 0 decimal place)

Exercise date = ((35*12*10)- (7*12*10))*32% =$1075.20 ($1075 if answer required to 0 decimal place)

Sale date = ((35*12*10)- (35*12*10))*32% = $0

Part B

Grant date = (7*12*10)*21% = $176.40 ($176 if answer required to 0 decimal place)

Exercise date = ((35*12*10)- (7*12*10))*21% =$705.60 ($706 if answer required to 0 decimal place)

Sale date = 176.40+705.60 = $882